OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5

Mehek (mehek14):

you need the y-intercept so far, you have \(y=\dfrac{3}{5}x+b\) b = y-intercept to solve for it, plug (2,-4) in it \(-4=\dfrac{3}{5}*2+b\\-4=1.2+b\\-5.2=b\)

Mehek (mehek14):

so your equation is \(y=\dfrac{3}{5}x-5.2\)

OpenStudy (anonymous):

what if they just give me this

OpenStudy (anonymous):

and the second box has an x already and and the fifth and sixth is Y =